OpenStudy (anonymous):

Solve algebraically then find the exact value of x: log x+log(x-2)=log 8

OpenStudy (anonymous):

start with \[\log(x(x-2))=\log(8)\] then since the log is a one to one function you know that \[x(x-2)=8\] and you can solve the resulting quadratic equation make sure to check your answers, because you cannot take the log of a negative number

OpenStudy (anonymous):

OK so x^2-2x=8 do you change that to x^2- 2x-8=0 and then factor it?